EXPERIMENT 6 CALORIMETRY II TABLE OF OBSERVATIONS AND CALCULATIONS Unchude the formulas and show all the detailed steps of the calculations). Part L- Heat Change of Endothermic and Exothermic Reactions; IA. Endothermie Reaction Observations Mass of aluminum cup+ stir bar+ (g) +stir bar (g water (g) Mass of al. cup Mass of NH4CI Tinitial, C Trial 1 Trial 2 Calculations Knowing that qsolution-4m, show the calculation of 솨lan for one of your trials. Trial 1 AH (kJ/g) Average ΔH (kJ/g):- 1B. Exothermic Reaction Observations Mass of Mass of al cup Mass of CaCTisida, CTiea, C aluminum cup stir bar+(g) + stir bar (g waer (g) 35.541| 135.20커 20s I a3.7 | Sa,S 3 5.59S 35.740 20.031 3.4 51.3 Trial 1 Trial 2 Calculations Knowing that qsolution--qnn, show the calculation of hn for one of your trials. Trial 1 AH (kJ/g) Average AH (kJ/g):
Answer #1

answering first question as multiple questions asked, which is against HomeworkLib QnA,

1. calculation of dHrxn for trial first,

given data

weight of Al = 35.53 , moles of Al = 35.53/27 = 1.32 moles

weight of Al+water = 135.85 so weight of water = 135.85-35.53 = 100.32 gm

and since q solution = - q rxn

heat capacity of water = 4.184 J/mol-K , specific heat capacity of NH4Cl = 84.1 J/mol-K, S for Al = 0.902

Delta H = m*s* delta T , m = mass of compound, s = specific heat of compound , delta T = 23.7-11.7 = 12 degree celcius

Put all above values in this equation,

delta H = (m1s1+m2s2+m3s3) * delta T

= (35.53*0.900 + 100.32*4.184 + 20.14*84.1) * 12 = 25.73 kJ/g

ANswer